# Discrete Fourier Transform versus Discrete Chi-Square Method

### v2

@@ Abstract
We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM). DCM is designed for detecting many signals superimposed on an unknown trend. The solution for the non-linear DCM model is an ill-posed problem. The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models. DCM is a simple numerical time series analysis method that performs a massive number of linear least squares fits. Hence, the data spacing, even or uneven, is irrelevant. We show that our numerical solution for the DCM model fulfils the three conditions of a well-posed problem: existence, uniqueness and stability. The Fisher-test is used to identify the best DCM model from all alternative tested DCM models. The correct DCM model must also pass our Predictivity-test. Our analyses of seven different simulated data samples expose the weaknesses of DFT and the efficiency of DCM. The DCM signal and trend detection depend only on the sample size and the accuracy of data. DCM is an ideal forecasting method because the time span of observations is irrelevant. If the Gauß-Markov theorem is valid, DCM can not fail. We recommend fast sampling of large high quality datasets and the analysis of those datasets using numerical DCM parallel computation Python code.
@@
Ill-posed problems, Time series analysis, Non-linear models, Discrete Chi-square Method (DCM), Discrete Fourier Transform (DFT)
@@ 1 Introduction
[21] formulated the least squares method. [11] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [12]. The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [30]. The Gauß-Markov theorem is the backbone of our Discrete Chi-Square Method (DCM). This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
@@
[…]
However, no published solution seems to exist for the more complicated cases, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines. We will show that DCM can find a well-posed numerical solution even for such complicated cases. This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend. The noise level can be unknown because the performance of DCM depends only on the sample size $(n)$ and data accuracy $(\sigma)$. The time span $(\Delta T)$ is irrelevant. The data spacing, even or uneven, is also irrelevant.
@@
We present the DCM formulation in Section 2.1. Data samples are simulated using the DCM model (Sect. 2.2). The four DFT limitations are discussed in Section 2.3 (Equations 26-29). DCM and DFT time series analysis of the simulated data samples shows that DCM does not suffer from the DFT limitations (Sects. 3.1-3.7). We then present the identification of the best DCM model for the data (Sect. 3.8), the DCM significance estimates (Sect. 3.9) and the DCM predictions (Sect. 3.10). Finally, we discuss our results (Sect. 4).
@@ 2.1 Discrete Chi-square method (DCM)
In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 19) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$.
@@
DCM determines the signal parameters
@@
We use the bootstrap technique to estimate the DCM model parameter errors [9]. The tested bootstrap frequencies are the same as in the short search. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create numerous artificial bootstrap data samples
@@
Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter. For each particular model parameter, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
@@ 2.2 Simulated DCM model data samples
We use seven different $g(t)$ models to simulate 21 artificial data samples (Sections 3.1-3.7). The time interval of all simulated samples is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [20] is always
@@ 2.3 Discrete Fourier Transform (DFT)
where $s_{\mathrm{DFT}}(t)$ is the pure sine signal and $p_{\mathrm{DFT}}(t)$ is the trend. The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively. Our DFT analysis may fail for four main reasons.
@@ 3 Results
We apply DCM and DFT to seven different simulated data samples (Sections 3.1 - 3.7). DCM analysis succeeds for all samples. DFT analysis fails for every sample.
@@ 3.7 Model 7
Our seventh simulation model is
@@
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$, $P_{1}=1.2$, $P_{2}=1.4$, $M_{0}=1$ $M_{1}=0.25$ and $M_{2}=0.5$. The coefficients $c_{1}=0.3687$, $c_{2}=0.7374$, $c_{3}=0.4000$, $c_{4}=0.3708$, $c_{5}=0.7416$ and $c_{6}=0.6000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{1,max,2}}$, $A_{2}$, $t_{\mathrm{2,min,1}}$, $t_{\mathrm{2,min,2}}$, $t_{\mathrm{2,max,1}}$ and $t_{\mathrm{2,max,2}}$ values given in Table 7 (Column 1). This simulated data sample is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 26). The parabolic $p(t)$ represents “a trend” (Equation 27). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” because $\Delta f=f_{1}-f_{2}=0.12<f_{0}=1$ (Equation 28). The two $h_{1}(t)$ and $h_{2}(t)$ signals are not “pure sines” (Equation 29). These signals are double waves having two unequal minima and maxima. All four main reasons that can cause the failure of DFT analysis are present (Equations 26-29). Our DCM and DFT time series analysis of Model 7 simulated data is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
The DCM orders of Model 7 are $K_{1}=2$, $K_{2}=2$ and $K_{3}=2$. This model has $\eta=13$ free parameters (Equation 8). We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 7 (Columns 2-4). The DCM analysis results are displayed for one sample of Model 7 simulated data (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$). Since there are $\eta=13$ free model parameters, this $n=100$ sample is quite small for time series analysis. The long and short DCM searches give $P_{1}=1.200$ and $P_{2}=1.385$ (Figure 7a: diamonds), and $P_{1}=1.199$ and $P_{2}=1.405$ (Figure 7b: Diamonds). The black $g(t)$ model line goes through all black $y_{i}$ data dots (Figure 7c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.950\pm 0.050$, $M_{1}=0.2447\pm 0.0067$ and $M_{2}=0.551\pm 0.050$ (Figure 7c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) are displayed in Figure 7d. The $n=100$ residuals (blue dots) are offset to the level of -1.8 (dotted blue line). These DCM model residuals are small and their level is stable. These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data samples $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
@@
The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 26-29.
@@ 3.8 Best model
Larger $F_{R}$ or $F_{\chi}$ values have smaller $Q_{F}$ critical levels. Hence, the probability for the $H_{0}$ hypothesis rejection increases when $F_{R}$ or $F_{\chi}$ increases. If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease. This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ (Equations 37 and 38: first terms). At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 37 and 38: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values. This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
@@ 3.9 Significance estimates
There is no need to discuss the DFT significance estimates [17, their Equation 22] because this method fails to detect the correct Model 1 period.
@@
Figure 8: DCM model prediction (Model 3 simulated data for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are predictive data). (a) One signal model $g_{1,1,0}$ results. Predictive data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and predicted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots). Continuous black line of predictive data model $g(t,\bm{\beta}_{\mathrm{pred}})$ ends to vertical dotted blue line, where dotted black line of predicted model $g^{\prime}(t,\bm{\beta}_{\mathrm{pred}})$ begins. Dotted red line denotes $\pm 3\sigma$ errors of both models. Residuals of predictive data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of predicted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line). (b) Two signal model $g_{2,1,0}$ results. Otherwise as in “a”. (c) Three signal model $g_{3,1,0}$ results. Otherwise as in “a”.
@@ 3.10 Prediction
There are numerous techniques for predicting (forecasting) a time series [e.g., 16, 18]. The DCM model $g(t_{i},\bm{\beta})$ can be used to predict. We divide all data into the predictive data and the predicted data. The time points, the observations and the errors of these samples are

### v3

# Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods

@@ Abstract
We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM). DCM is designed for detecting signal(-s) superimposed on an unknown trend. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational statistical well-posed solution for this problem. The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models. DCM can not fail because this simple time series analysis method computes a massive number of linear least squares fits. Hence, the data spacing, even or uneven, is irrelevant. We use the Fisher-test to identify the best DCM model from all alternative tested DCM models. This best model must also pass our Predictivity-test. Our analyses of simulated complex data sets expose the weaknesses of DFT and the efficiency of DCM. The DFT and DCM are frequency-domain parametric methods. There are many other similar parametric methods. Just like DFT, those other methods suffer from their own particular application limitations. The list of those limitations is long. DCM suffers from none of those limitations. The performance of DCM depends only on the data set size and accuracy. DCM is an ideal forecasting method because the data set time span $(\Delta T)$ is irrelevant. It does not matter how long $(\Delta T)$ and/or complex the data set is because DCM will inevitably detect the signal(-s) and the trend when the data set size $(n)$ and/or accuracy $(\sigma)$ become adequate.
@@
Computational statistics, Ill-posed problems, Non-linear time series analysis models, Discrete Chi-square Method, Discrete Fourier Transform
@@ 1 Introduction
[24] formulated the least squares method. [13] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [14]. The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [34]. The Gauß-Markov theorem is the backbone of our DCM. This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
@@
Ill-posed problems arise in many fields of science [27, 30, 25]. Typical examples are partial differential equation solutions [17], nonlinear parameter estimation [1] and regularisation of inverse problems [9, 33]. Ill-posed problems are encountered in time series analysis [3, 5] because the models for the data are often non-linear [32, 31]. The unknown free parameters of the model, namely the frequencies, are in the arguments of trigonometric functions representing the signals. This causes non-linearity because the model partial derivatives contain free parameters of the model.
There are numerous techniques for solving the ill-posed problem of detecting one signal superimposed on a trend, especially for evenly spaced data [7, 28, 2]. Kay and Marple [22] compared many spectral estimating techniques that can detect more than one signal from evenly spaced data. They discussed how the time span of data $(\Delta T)$ limits the frequency resolution for detecting many signals and causes leakage that shifts power from one spectral peak to another. The spectral estimating techniques fail if the data are non-stationary. Therefore, trends changing the sample mean and/or variance must be removed before applying spectral estimation [26]. Ghaderpour et al. [15] presented different techniques for reducing spectral leakage. They also discussed techniques for detecting many signals in unevenly spaced data, if these signals are pure sines.
However, no published solution seems to exist for the more complex data sets, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines. We will show that DCM can find a well-posed computational statistical solution even for such complex data sets. This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend. The noise level can be unknown because the performance of DCM depends only on the sample size $(n)$ and data accuracy $(\sigma)$. The time span $(\Delta T)$ is irrelevant. The data spacing, even or uneven, is also irrelevant.
@@
We present the DCM formulation in Section 2.1. Data sets are simulated using the DCM model (Sect. 2.2). The four DFT application limitations are discussed in Section 2.3 (Equations 27-30). DCM and DFT time series analysis of the simulated data sets shows that DCM does not suffer from the DFT application limitations (Sects. 3.1-3.7). We then present the identification of the best DCM model for the data (Sect. 3.8), the DCM significance estimates (Sect. 3.9) and the DCM predictions (Sect. 3.10). Finally, we discuss our results (Sect. 4).
@@ 2.1 Discrete Chi-square method (DCM)
In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 19) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$. These one-dimensional periodogram slices (Equation 20) allow us "to see inside” the multi-dimensional structure of the DCM test statistic $z$ (Equations 15 and 16). This kind of visualisation is important in time series analysis [29].
@@
Furlan and Mortarino [11] emphasise that the analytical error estimates for the non-linear model free parameters can be tricky. They compare different analytical free parameter error estimating methods by using the computational statistical bootstrap technique [10]. They conclude that the analytical error estimates become less reliable if the number of free parameters increases. For our non-linear DCM model, we determine the $i$:th signal parameters
@@
Of these model parameters, all $P_{i}=1/f_{i}$ and all $M_{k}$ estimates are among the free parameters $\bm{\beta}$. However, the $A_{i}$, $t_{\mathrm{i,min,1}}$, $t_{\mathrm{i,min,2}}$, $t_{\mathrm{i,max,1}}$ and $t_{\mathrm{i,max,2}}$ model parameters depend on $\bm{\beta}$ values. The analytical solutions for these model parameters are simple for pure sines $(K_{2}=1)$, but become quite complicated for double waves $(K_{2}=2)$. Clearly, the analytical solution for the errors of DCM model parameters would be tedious. The analytical solution for the model error
$g(t)\pm\sigma_{g(t)}$ (22)
would be practically impossible because the number of DCM model free parameters is large, especially if the data contains many signals. We use the computational statistical bootstrap technique [10] to solve these analytical problems. This approach gives the DCM model parameter error estimates, as well as the $\sigma_{g(t)}$ model error estimates (Equation 22). In our bootstrap, the tested frequencies are the same as in the short search. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create numerous artificial bootstrap data sets
@@
Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter value and model $g(t)$ value. For each particular model parameter and each particular $g(t)$, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
@@ 2.2 Simulated DCM model data samples
We use seven different $g(t)$ models to simulate 21 artificial complex data sets (Sections 3.1-3.7). The time interval of all simulated data sets is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [22] is always
@@ 2.3 Discrete Fourier Transform (DFT)
where $s_{\mathrm{DFT}}(t)$ is the pure sine signal and $p_{\mathrm{DFT}}(t)$ is the trend. The pure sine signals for the $y_{i}$ data and the $\epsilon_{i}$ residuals are denoted with $s_{y,\mathrm{DFT}}(t)$ and $s_{\epsilon,\mathrm{DFT}}(t)$, respectively. Our DFT analysis may fail due to four application limitations.
@@ 3 Results
We apply DCM and DFT to seven different simulated data sets (Sections 3.1 - 3.7). As we proceed, these data sets become increasingly complex. DCM analysis succeeds for all data sets. DFT analysis fails for every data set.
@@ 3.7 Model 7
In this section, we analyse our most complex data set. The seventh simulation model is
@@
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$, $P_{1}=1.2$, $P_{2}=1.4$, $M_{0}=1$ $M_{1}=0.25$ and $M_{2}=0.5$. The coefficients $c_{1}=0.3687$, $c_{2}=0.7374$, $c_{3}=0.4000$, $c_{4}=0.3708$, $c_{5}=0.7416$ and $c_{6}=0.6000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{1,max,2}}$, $A_{2}$, $t_{\mathrm{2,min,1}}$, $t_{\mathrm{2,min,2}}$, $t_{\mathrm{2,max,1}}$ and $t_{\mathrm{2,max,2}}$ values given in Table 7 (Column 1). This simulated data sample is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 27). The parabolic $p(t)$ represents “a trend” (Equation 28). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” because $\Delta f=f_{1}-f_{2}=0.12<f_{0}=1$ (Equation 29). The two $h_{1}(t)$ and $h_{2}(t)$ signals are not “pure sines” (Equation 30). These signals are double waves having two unequal minima and maxima. All four main reasons that can cause the failure of DFT analysis are present (Equations 27-30). Therefore, this simulated data set is the most complex one of all analysed seven data sets. Our DCM and DFT time series analysis of Model 7 simulated data is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
The DCM orders of Model 7 are $K_{1}=2$, $K_{2}=2$ and $K_{3}=2$. This model has $\eta=13$ free parameters (Equation 8). We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 7 (Columns 2-4). The DCM analysis results are displayed for one sample of Model 7 simulated data (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$). Since there are $\eta=13$ free model parameters, this $n=100$ sample is quite small for time series analysis. The long and short DCM searches give $P_{1}=1.200$ and $P_{2}=1.385$ (Figure 7a: diamonds), and $P_{1}=1.199$ and $P_{2}=1.405$ (Figure 7b: Diamonds). The black $g(t)$ model line goes through all black $y_{i}$ data dots (Figure 7c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.950\pm 0.050$, $M_{1}=0.2447\pm 0.0067$ and $M_{2}=0.551\pm 0.050$ (Figure 7c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) are displayed in Figure 7d. The $n=100$ residuals (blue dots) are offset to the level of -1.8 (dotted blue line). These DCM model residuals are small and their level is stable. These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data sets $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
@@
The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 27-30.
@@ 3.8 Best model
Larger $F_{R}$ or $F_{\chi}$ values have smaller $Q_{F}$ critical levels. Hence, the probability for the $H_{0}$ hypothesis rejection increases when $F_{R}$ or $F_{\chi}$ increases. If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease. This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ increase (Equations 38 and 39: first terms). At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 38 and 39: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values. This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
@@ 3.9 Significance estimates
There is no need to discuss the DFT significance estimates [19, their Equation 22] because this method fails to detect the correct Model 1 period.
@@
Figure 8: DCM model prediction (Model 3 simulated data for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are predictive data). (a) One signal model $g_{1,1,0}$ results. Predictive data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and predicted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots). Continuous black line of predictive data model $g(t,\bm{\beta}_{\mathrm{pred}})$ ends to vertical dotted blue line, where dotted black line of predicted model $g^{\prime}(t,\bm{\beta}_{\mathrm{pred}})$ begins. Dotted red line denotes $\pm 3\sigma_{g(t)}$ errors of both models (Equation 22). Residuals of predictive data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of predicted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line). (b) Two signal model $g_{2,1,0}$ results. Otherwise as in “a”. (c) Three signal model $g_{3,1,0}$ results. Otherwise as in “a”.
@@ 3.10 Prediction
There are numerous techniques for predicting (forecasting) a time series [e.g., 18, 20, 23]. The DCM model $g(t_{i},\bm{\beta})$ can be used to predict. We divide all data into the predictive data and the predicted data. The time points, the observations and the errors of these samples are
@@
The analyses of all seven complex data sets indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7). This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [22, 15] would explain why correct signal periods have not been detected earlier from the sunspot data.